given Investment = 900, consumption=600+0.6yd
calculate the the equilibrium level of income
1
Expert's answer
2012-08-14T10:59:49-0400
Y = C+I
Y= 600 + 0.6Y + 900
0.4Y = 1500
Y = 1500/0.4
Y=3750